OpenStudy (anonymous):

(1/3)x^(-2/3) - (1/5)x^(-4/5) = 0

OpenStudy (anonymous):

might try to factor out x^(-4/5) to begin?

OpenStudy (jdoe0001):

\(\large \bf \cfrac{1}{3}x^{-\frac{2}{3}}-\cfrac{1}{5}x^{-\frac{4}{5}}=0\quad ?\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

came from taking the derivative of a function

OpenStudy (jdoe0001):

k

OpenStudy (anonymous):

got to a point where the equation is x= (125/27) * x^(15/4) but may have gone wrong

OpenStudy (jdoe0001):

\(\Large { \cfrac{1}{3}x^{-\frac{2}{3}}-\cfrac{1}{5}x^{-\frac{4}{5}}=0\implies \cfrac{x^{-\frac{2}{3}}}{3}-\cfrac{x^{-\frac{4}{5}}}{5}=0 \\ \quad \\ \cfrac{x^{-\frac{2}{3}}}{3}=\cfrac{x^{-\frac{4}{5}}}{5}\implies \cfrac{x^{-\frac{2}{3}}}{x^{-\frac{4}{5}}}=\cfrac{3}{5}\implies x^{-\frac{2}{3}+\frac{4}{5}}=\cfrac{3}{5} \\ \quad \\ x^{\frac{2}{15}}=\cfrac{3}{5}\implies \left(x^{\cancel{ \frac{2}{15}}}\right)^{\cancel{ \frac{15}{2}}}=\left(\cfrac{3}{5}\right)^{\frac{15}{2}} }\) thus far it looks like

OpenStudy (anonymous):

k

OpenStudy (anonymous):

would that be plus/minus, since we took the square root of a square at some point?

OpenStudy (jdoe0001):

hmm well... can be either I gather... we started off with a rational exponent though but since is square root... could be either, depending on context